OpenStudy (magbak):

Medal will be awarded!!!! Part 1: Create your own factorable trinomial. (1 point) Part 2: Explain, in complete sentences, how the trinomial is factored. (1 point) Part 3: Explain, in complete sentences, the process used to check the factors for accuracy. (1 point)

OpenStudy (mertsj):

Write a binomial and post it.

OpenStudy (magbak):

x^2 + 2

OpenStudy (mertsj):

Write another binomial and post it.

OpenStudy (magbak):

x + 4

OpenStudy (mertsj):

Ok. Make the first on x+2. It will make it easier on yourself.

OpenStudy (mertsj):

Now multiply those two binomials.

OpenStudy (texaschic101):

Opposite of FOIL can factor trinomials...right ?

OpenStudy (mertsj):

Ok. texas wants to help you now. Too many cooks, you know....

OpenStudy (magbak):

Ok but one thing is I need the trinomial to have a squared x value like x^2 + 2x + 3

OpenStudy (texaschic101):

I do not want to take over this question....@Mertsj ...its yours

OpenStudy (mertsj):

If you multiply the two binomials you will have a trinomial and it will have an x^2 in it.

OpenStudy (magbak):

Ok.

OpenStudy (magbak):

so the factors are (x +2) (x + 2)

OpenStudy (mertsj):

What are the two binomials you are going to multiply?

OpenStudy (magbak):

(x+2 ) (x+2)

OpenStudy (mertsj):

When I said write a binomial you wrote x^2+2. I said to make it x+2. Then I said write another binomial and you wrote x+4 Then I said to multiply the two binomials and then you got off on this (x+2)(x+2) stuff somehow.

OpenStudy (mertsj):

So: Multiply (x+2)(x+4) then you will have a trinomial and it will have an x^2 in it

OpenStudy (primeralph):

...........Suddenly, a wild cat appears. _ヽ   \\ Λ_Λ    \( ˇωˇ)      > ⌒ヽ    /   へ\    /  / \\    レ ノ   ヽ_つ   / /   / /|  ( (ヽ  | |、\  | 丿 \ ⌒)  | |  ) / `ノ )  Lノ (_/.

OpenStudy (magbak):

Ok I am sorry but I did not understand that part. So (x+2)(x+4) is x^2 + 6x + 8

OpenStudy (mertsj):

Yes. Now factor x^2+6x+8

OpenStudy (primeralph):

Uh I think there's a disconnection somewhere. Maybe you should start again.

OpenStudy (mertsj):

You have now done part 1 which says to create your own factorable trinomial.

OpenStudy (mertsj):

Now you are on part 2 which says to factor your trinomial.

OpenStudy (magbak):

It is (x+2)(x+4)

OpenStudy (mertsj):

That is correct.

OpenStudy (magbak):

Now what do I do.

OpenStudy (mertsj):

Now do part 3. How would you know that the factors are correct? Hint: Are 3 and 4 factors of 12? How do you know?

OpenStudy (primeralph):

.........but it's a trinomial. Shouldn't it be 3 roots?

OpenStudy (magbak):

You just multiply the factors again and if it gives you the trinomial the factors are correct. RIGHT???

OpenStudy (mertsj):

yes. That is right

OpenStudy (texaschic101):

thats what I would do :)

OpenStudy (magbak):

Thank you all.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

TRINOMIAL: \[x ^{2}+5x+6\] Fatorization: \[x^{2}+5x+6\]=\[x ^{2}+2x+3x+6=x(x+2)+3(x+2)=(x+3)(x+2)\] first break the number with x in such away that the product of the numbers after breaking it is the number to itz right here5x=2x+3x and 2.3=6 part 3: multiply both the expressions i.e.,\[(x+3)(x+2)=x^{2}+5x+6\]

OpenStudy (magbak):

WOW WAY TO THROW ME OF ?!?!?!?!?!?!?!

OpenStudy (texaschic101):

lol....it gets confusing sometimes

OpenStudy (magbak):

Yea I can see that but why was the other mans explanation much easier.

OpenStudy (texaschic101):

stick with the one you are more comfortable with